OpenStudy (anonymous):

A store is having a sale on trail mix and jelly beans. For 2 pounds of trail mix and 3 pounds of jelly beans, the total cost is $12 . For 6 pounds of trail mix and 5 pounds of jelly beans, the total cost is $30 . Find the cost for each pound of trail mix and each pound of jelly beans.

OpenStudy (anonymous):

let the cost of a pound of trail mix be x and the cost of a pound of jelly bean be y 2x +3y = 12 and 6x+5y = 30

OpenStudy (anonymous):

i get y = -2/3x+12

OpenStudy (anonymous):

y= -6/5x+30

OpenStudy (jhannybean):

hint: you don't necessarily have to solve for y

OpenStudy (anonymous):

so whats the answer :o ?

OpenStudy (jhannybean):

I think the process of elimination would work better

OpenStudy (anonymous):

how do you do process of elimination on this ?

OpenStudy (anonymous):

let x be a pound of trail mix and let y be a pound of jelly beans. 2x+3y=12 6x+5y=30 Let's do substitution. Isolate x in the first equation. 2x= 12-3y x=6-3/2 y. Substitute that into the second equation. 6(6-3/2y)+5y=30. 36-18/2 y+5y=30 -9y+5y=-6 -4y=6 y=1.5 Put that back into the first equation 2x=12-3(1.5) 2x=7.5 x=3.75 So the trail mix is $3.75 and the jellybeans are $1.50

OpenStudy (jhannybean):

1. 2x +3y = 12 2. 6x+5y = 30 multiply 1. by -3 to eliminate the x -6x -9y = -36 6x +5y = 30

OpenStudy (jhannybean):

Then you add these two equations together. (-9 +5)y = -36+30 -4y = -6 y = 6/4 = 3/2

OpenStudy (anonymous):

oh okay i get it now !!

OpenStudy (jhannybean):

Then you use y = 3/2 into any one of the two equations to find x. 2x + 3(3/2) = 12 2x + 9/2 = 12 2x = 12 - 9/2 2x = 15/2 x = (15/2)/2 = 15/4